# Mapping Overlaps in Benchmarks through Perplexity in the Wild

### v1

@@ 1 Introduction
Perplexity provides a useful lens for quantifying relationships between pretraining, skills, and benchmark performance. Low perplexity on a passage suggests that the model has seen similar linguistic and conceptual patterns during training and is familiar with the content. High perplexity, by contrast, indicates unfamiliarity and underrepresentation in pretraining. Thus, the distribution of perplexity values across large corpora serves as a fingerprint of the model’s training exposure and more or less acquired capacity. Importantly, because different benchmarks stress different capabilities, they map onto different perplexity distributions when probing across the same corpus. In other words, corpora encode benchmark signatures because benchmarks are not foreign entities imposed on the model after training, but rather structured samplings of capabilities that themselves emerge from the distribution of in-the-wild data. Perplexity serves as the bridge between corpus exposure and benchmark performance, making it possible to identify and characterize these signatures without requiring direct evaluation on the benchmark itself  [^2] More discussions about prior works see A.1.. We therefore leverage perplexity as the basis and covariate for salient token selection and signature formation  [^3] More details see Section 3 and Appendix A.2.. The following three levels in this work provide a holistic framework: semantics address task design, performance captures model behavior, and signatures reveal a fingerprint of model capacity. The overlap between benchmarks across each of these levels highlights the interconnected capacity space - an oft-discussed yet difficult-to-formalize concept and so represents a promising tool for evaluating benchmark validity. This rationale is illustrated in Figure 2.
@@ 2 Semantic Overlaps and Performance Overlaps
Overall, this mitigates sample-size bias and yields a more robust similarity estimate. Full procedural details appear in Appendix A.4.1. The overlap between $B_{a}$ and $B_{b}$ is defined by the $\widehat{A}_{\mathrm{sem}}(B_{a},B_{b})$.
@@ 3 Mining Benchmark Signatures from In-the-wild data
The overall process of mining signatures can be found in Algorithm 1 and its details can be found in Appendix 4. Let $d$ denote the number of “in-the-wild” tokens (in our case, tokens drawn from large-scale pretraining corpora [^4] Progressing from fine to coarse granularity, we have token-, chunk-, and document-level perplexities. We provide more experimental results of why the token-level operation is the best. Details are shown in Appendix A.2.), denoted as $\mathcal{T}=\{t_{1},\dots,t_{d}\}$, where $d$ typically scales to billions. For any benchmark $B_{j}$, our objective in extracting its benchmark signature is to isolate a subset of salient tokens $\mathcal{T}_{j}^{\prime}=\{t^{\prime}_{1},\dots,t^{\prime}_{d^{\prime}}\}\subset\mathcal{T}$ that are maximally informative in explaining variations in LLM performance. We formalize this as a regression problem: let $\mathbf{\hat{y}}_{j}:=(y_{1,j},\dots,y_{m,j})^{\top}\in\mathbb{R}^{m}$ denote the performance of $m$ language models on benchmark $B_{j}$. The covariate matrix $\mathbf{P}\in\mathbb{R}^{m\times d}$ contains token-level perplexities, where entry $\mathbf{P}_{ij}\equiv p_{ij}$ corresponds to the perplexity of token $t_{j}$ under LLM $M_{i}$. The challenge lies in the high-dimensional regime ($d\gg m$), where classical regression approaches are ill-posed. To make progress, we must uncover and exploit latent structural properties of the problem. In particular, we put forward a key assumption and a follow-up question:
@@ 3.1 Token-level Filtering with Perplexity Correlations
    Justification from the Ultra-High Dimensional Regime: Our problem, with $d\gg m$, resides in the ultra-high dimensional setting. Theoretical frameworks developed for this regime, such as Sure Independence Screening (SIS; Fan & Lv (2008)), provide formal guarantees for marginal screening. The “sure screening property” ensures that, under sparsity and certain regularity conditions, correlation-based filtering can discard the vast majority of irrelevant features while retaining the true predictive signals with very high probability. We further explain how several key conditions of SIS are plausible in our context in Section A.3.
@@ 3.3 Signature-Level Overlap
Consider two benchmark signature vectors, $S_{1}$ and $S_{2}$, each including several pieces of context (30 pieces separated by space) + the salient token. We use 32 models to process these signatures, reading their respective pre-contexts, producing the last token-level perplexities and calculating overlaps. If the models are confused to a similar degree by both signatures, that is a strong indicator that the two benchmarks align. Since some “weak” models consistently produce high perplexity, we normalize each model’s perplexity into its z-score within the model. We then compute the mean of z-scored perplexities of the two benchmark signatures within each model and the Spearman correlation between these two mean lists to represent the signature-level overlap, aligning with performance level results and indicating models’ relative relation of perplexity and skill familiarity on the signature. Refer to Appendix A.5.1 for a formalized walk-through.
@@ 4 Results
Our experiments are conducted on 32 models and 88 benchmarks, including many of the most widely used ones. We extract benchmark signatures from the open dataset RedPajama (Weber et al., 2024). See Appendix A.5 for the full details of the experimental setup.
@@ 5 Final Remarks
Appendix A Appendix
A.1 Related Literature
@@
A.2 Comparisons between token-, chunk-, and document-level perplexity
@@
A.3 Conditions for Sure Independence Screening (SIS)
@@
A.4 Technical Details
A.4.1 Semantic-level Bootstrapped Similarity Calculation
@@
A.4.2 AIC Stepwise Forward Selection Algorithm
@@ A.5 Experiment Setup
A.5.1 Experiment Walkthrough
As discussed, we measure perplexity at three granularities - token, chunk, and document levels (from fine to coarse). The segmentation procedure for each is detailed in §A.2. We ultimately focus on the token level because it provides the clearest view of prominent signals for the pre-filtering stage.
@@ A.5.1 Experiment Walkthrough
Using the vLLM setup described in §A.5.4, we evaluate 32 models on the token contexts and extract token-level perplexities, forming the covariate (feature) matrix
@@
In parallel, we compute model performance on a series of benchmarks and subfields using the lm-evaluation-harness (details in §A.5.5). Let
@@ A.5.2 Datasets
A.5.3 Models Used
The models we used in this study are summarized in Table 4. Each entry cites the official paper if available, otherwise the model card.
Model (HF Repository) Citation
meta-llama/Llama-3.1-8B-Instruct Grattafiori et al. (2024)
meta-llama/Llama-3.2-1B-Instruct Grattafiori et al. (2024)
meta-llama/Llama-3.2-3B-Instruct Grattafiori et al. (2024)
google/gemma-3-4b-it Team (2025a)
google/gemma-3-12b-it Team (2025a)
google/gemma-3-27b-it Team (2025a)
mistralai/Mistral-7B-Instruct-v0.3 Jiang et al. (2023)
deepseek-ai/deepseek-llm-7b-chat DeepSeek-AI et al. (2024)
Qwen/Qwen3-0.6B Team (2025b)
Qwen/Qwen3-1.7B Team (2025b)
Qwen/Qwen3-4B Team (2025b)
Qwen/Qwen3-8B Team (2025b)
tiiuae/falcon-rw-1b Penedo et al. (2023)
EleutherAI/pythia-1b Biderman et al. (2023)
EleutherAI/pythia-6.9b-v0 Biderman et al. (2023)
EleutherAI/pythia-12b-deduped Biderman et al. (2023)
mosaicml/mpt-7b-instruct Team (2023)
microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024a)
microsoft/Phi-4-mini-instruct Microsoft et al. (2025)
microsoft/Phi-4 Abdin et al. (2024b)
01-ai/Yi-1.5-9B-Chat AI et al. (2025)
01-ai/Yi-1.5-6B-Chat AI et al. (2025)
mistralai/Ministral-8B-Instruct-2410 Jiang et al. (2024)
openai-community/gpt2-medium Radford et al. (2019)
openai-community/gpt2-large Radford et al. (2019)
openai-community/gpt2-xl Radford et al. (2019)
zai-org/chatglm3-6b GLM et al. (2024)
zai-org/glm-4-9b-chat-hf GLM et al. (2024)
zai-org/codegeex4-all-9b Zheng et al. (2023)
allenai/OLMo-2-1124-13B-Instruct OLMo et al. (2024)
allenai/OLMo-2-1124-7B-Instruct OLMo et al. (2024)
allenai/OLMo-2-0425-1B-Instruct OLMo et al. (2024)
Table 4: Models used in this study. Multiple variants may share the same citation.
A.5.4 Inference with vLLM

### v2

@@ 1 Introduction
Perplexity provides a useful lens for quantifying relationships between pretraining, skills, and benchmark performance. Low perplexity on a passage suggests that the model has seen similar linguistic and conceptual patterns during training and is familiar with the content. High perplexity, by contrast, indicates unfamiliarity and underrepresentation in pretraining. Thus, the distribution of perplexity values across large corpora serves as a fingerprint of the model’s training exposure and more or less acquired capacity. Importantly, because different benchmarks stress different capabilities, they map onto different perplexity distributions when probing across the same corpus. In other words, corpora encode benchmark signatures because benchmarks are not foreign entities imposed on the model after training, but rather structured samplings of capabilities that themselves emerge from the distribution of in-the-wild data. Perplexity serves as the bridge between corpus exposure and benchmark performance, making it possible to identify and characterize these signatures without requiring direct evaluation on the benchmark itself  [^2] More discussions about prior works see A.. We therefore leverage perplexity as the basis and covariate for salient token selection and signature formation  [^3] More details see Section 3 and Appendix B.. The following three levels in this work provide a holistic framework: semantics address task design, performance captures model behavior, and signatures reveal a fingerprint of model capacity. The overlap between benchmarks across each of these levels highlights the interconnected capacity space - an oft-discussed yet difficult-to-formalize concept and so represents a promising tool for evaluating benchmark validity. This rationale is illustrated in Figure 2.
@@ 2 Semantic Overlaps and Performance Overlaps
Overall, this mitigates sample-size bias and yields a more robust similarity estimate. Full procedural details appear in Appendix D.1. The overlap between $B_{a}$ and $B_{b}$ is defined by the $\widehat{A}_{\mathrm{sem}}(B_{a},B_{b})$.
@@ 3 Mining Benchmark Signatures from In-the-Wild Data
The overall process of mining signatures can be found in Algorithm 1 and its details can be found in Appendix 4. Let $d$ denote the number of “in-the-wild” tokens (in our case, tokens drawn from large-scale pretraining corpora [^4] Progressing from fine to coarse granularity, we have token-, chunk-, and document-level perplexities. We provide more experimental results of why the token-level operation is the best. Details are shown in Appendix B.), denoted as $\mathcal{T}=\{t_{1},\dots,t_{d}\}$, where $d$ typically scales to billions. For any benchmark $B_{j}$, our objective in extracting its benchmark signature is to isolate a subset of salient tokens $\mathcal{T}_{j}^{\prime}=\{t^{\prime}_{1},\dots,t^{\prime}_{d^{\prime}}\}\subset\mathcal{T}$ that are maximally informative in explaining variations in LLM performance. We formalize this as a regression problem: let $\mathbf{\hat{y}}_{j}:=(y_{1,j},\dots,y_{m,j})^{\top}\in\mathbb{R}^{m}$ denote the performance of $m$ language models on benchmark $B_{j}$. The covariate matrix $\mathbf{P}\in\mathbb{R}^{m\times d}$ contains token-level perplexities, where entry $\mathbf{P}_{ij}\equiv p_{ij}$ corresponds to the perplexity of token $t_{j}$ under LLM $M_{i}$. The challenge lies in the high-dimensional regime ($d\gg m$), where classical regression approaches are ill-posed. To make progress, we must uncover and exploit latent structural properties of the problem. In particular, we put forward a key assumption and a follow-up question:
@@ 3.1 Token-level Filtering with Perplexity Correlations
    Justification from the Ultra-High Dimensional Regime: Our problem, with $d\gg m$, resides in the ultra-high dimensional setting. Theoretical frameworks developed for this regime, such as Sure Independence Screening (SIS; Fan & Lv (2008)), provide formal guarantees for marginal screening. The “sure screening property” ensures that, under sparsity and certain regularity conditions, correlation-based filtering can discard the vast majority of irrelevant features while retaining the true predictive signals with very high probability. We further explain how several key conditions of SIS are plausible in our context in Appendix C.
@@ 3.3 Signature-Level Overlap
Consider two benchmark signature vectors, $S_{1}$ and $S_{2}$, each including several pieces of context (30 pieces separated by space) + the salient token. We use 32 models to process these signatures, reading their respective pre-contexts, producing the last token-level perplexities and calculating overlaps. If the models are confused to a similar degree by both signatures, that is a strong indicator that the two benchmarks align. Since some “weak” models consistently produce high perplexity, we normalize each model’s perplexity into its z-score within the model. We then compute the mean of z-scored perplexities of the two benchmark signatures within each model and the Spearman correlation between these two mean lists to represent the signature-level overlap, aligning with performance level results and indicating models’ relative relation of perplexity and skill familiarity on the signature. Refer to Appendix E.1 for a formalized walk-through.
@@ 4 Results
Our experiments are conducted on 32 models and 88 benchmarks, including many of the most widely used ones. We extract benchmark signatures from the open dataset RedPajama (Weber et al., 2024). See Appendix E for the full details of the experimental setup.
@@ 5 Final Remarks
Appendix
Appendix A Related Literature
@@
Appendix B Comparisons between Token-, Chunk-, and Document-level Perplexity
@@
Appendix C Conditions for Sure Independence Screening (SIS)
@@
Appendix D Technical Details
D.1 Semantic-level Bootstrapped Similarity Calculation
@@
D.2 AIC Stepwise Forward Selection Algorithm
@@ Appendix E Experiment Setup
E.1 Experiment Walkthrough
As discussed, we measure perplexity at three granularities - token, chunk, and document levels (from fine to coarse). The segmentation procedure for each is detailed in §B. We ultimately focus on the token level because it provides the clearest view of prominent signals for the pre-filtering stage.
@@ E.1 Experiment Walkthrough
Using the vLLM setup described in §E.3, we evaluate 32 models on the token contexts and extract token-level perplexities, forming the covariate (feature) matrix
@@
In parallel, we compute model performance on a series of benchmarks and subfields using the lm-evaluation-harness (details in §E.4). Let
@@ E.2 Datasets
E.3 Inference with vLLM